# Cross-Modal Reconstruction Pretraining for Ramp Flow Prediction at Highway Interchanges

### v1

@@ 2.3 Pretraining framework for traffic prediction
5 Model Evaluation: Validation method used (VM), including real-world (RW) or simulation/benchmark (SB) data, Real-Time Capability (RTC), and Scenario Generalization Ability (SGA).
Abbreviations: RFP: Ramp Flow Prediction, MFP: Mainline Flow Prediction, NFP: Network-Wide Flow Prediction,
M-S: Mainline Short time series, M-LS: Mainline Long and Short time series, R-S: Ramp Short time series,
M-C: Mainline Complete data, R-C: Ramp Complete data, R: Robust to missing data,
RT: Real-Time, Pre-N: pretraining on Non-real-time data, Test-RT: Test on Real-Time data,
T: Traffic Flow Model, ML: Machine Learning, TL: Transfer Learning, Pre: Pretraining Model.
Table 1 summarizes representative methods across several dimensions, including data conditions, modeling paradigms, dependency modeling, and generalization ability. The comparison shows that pretraining frameworks are more effective in handling missing data and capturing complex spatiotemporal dependencies. This provides a solid foundation for the method proposed in this study.
@@ 3 Methodology
This section describes the proposed model in detail. The overall framework of the model is illustrated in Figure 1. The STDAE model consists of two components: a pretraining model and a downstream prediction model. In the pretraining stage, to more efficiently capture the spatio-temporal dependencies of traffic data, we design an STDAE structure. This structure comprises two mutually independent modules: a Temporal AutoEncoder (TAE) and a Spatial AutoEncoder (SAE). Both modules share the same architecture but focus separately on reconstructing features in the temporal and spatial dimensions, respectively, to generate temporal and spatial representations of ramp traffic flow states. In the downstream prediction stage, these representations are processed through a multi-layer perceptron (MLP) module and fused with the hidden layer of the prediction model to obtain the final forecast of ramp traffic flow.
@@ 3.1 Problem Statement
In this study, the short-term prediction of ramp traffic flow at interchanges is defined as a temporal prediction problem. Consider a hub interchange with a double-line cross, where the mainline consists of $N=8$ directions, including four upstream and four downstream directions for the two intersecting lines. There are $M=12$ traffic movements (ramps), with each ramp connecting the upstream and downstream mainline flows from two corresponding directions.
@@
To model the dependencies among ramps, we construct a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where each node represents a ramp, and edges represent spatial or functional correlations between ramps. The graph has $M$ nodes corresponding to the $M$ ramps, and its adjacency matrix $\mathbf{A}\in\mathbb{R}^{M\times M}$ encodes the pairwise relationships among ramps:
@@
In this study, we assume a fully connected graph by default, i.e., every ramp node is connected to all other ramp nodes, allowing the model to capture interactions among all ramps. This $M\times M$ fully connected graph structure is critical for accurate ramp traffic flow prediction.
3.2 The pretraining model
During the pretraining stage, we do not adopt a conventional autoencoding scheme that encodes the input into hidden representations and then reconstructs the input itself. Instead, the model reconstructs the ramp traffic flow sequences by leveraging the long historical time series of upstream and downstream mainline traffic flows. Spatiotemporal decoupling is the key strategy for modeling the spatiotemporal relationships between mainline and ramp traffic flows. Its core idea is to separate spatial dependencies from temporal dependencies and extract their features independently, thereby capturing complex spatiotemporal dynamics more accurately. In addition, an optional masking module is incorporated to address cases where part of the mainline traffic flow data is missing.
Figure 1: STDAE Pretraining and Prediction Framework
@@ 3.2.1 Input Feature Encoding
The input feature encoding is divided into two parts, namely feature fusion and feature encoding. In the feature fusion stage, we combine the mainline attributes—including traffic flow, speed, time of day, and day of week—represented as $\mathbf{V}_{t-(T_{\text{long}}-1):t}\in\mathbb{R}^{T_{\text{long}}\times N\times F}$. These features are reorganized into $\mathbf{X}\in\mathbb{R}^{T_{\text{long}}\times M\times C}$, where the feature representation of each ramp is constructed by aggregating the features from the corresponding upstream and downstream mainline connections. It is worth noting that this feature fusion step is required in both the pretraining and prediction stages, ensuring consistent representation of ramp-related inputs throughout the entire modeling process.
In the feature encoding stage, the input feature tensor $\mathbf{X}\in\mathbb{R}^{T_{\text{long}}\times M\times C}$ is selectively processed through a temporal mask or a spatial mask. If the mainline traffic data is complete, no masking is applied. When the task involves missing mainline data in certain directions or time intervals, the corresponding mask is used to simulate these missing values. To reduce computational cost for long sequences $T_{\mathrm{long}}$, the sequence is divided into $P$ non-overlapping segments, each of length $L=T_{\mathrm{long}}/P$, yielding $\mathbf{Z}\in\mathbb{R}^{M\times P\times LC}$. A fully connected layer then maps $\mathbf{Z}$ to an embedding $\mathbf{E}_{p}\in\mathbb{R}^{M\times P\times D}$. It is first passed through a layer normalization and then added to a 2D sinusoidal positional encoding $\mathbf{E}_{\mathrm{pos}}$ (Wang and Liu, 2021) to obtain the final input embedding $\mathbf{E}\in\mathbb{R}^{M\times P\times D}$.
@@ 3.2.2 STDAE
The embedding $\mathbf{E}$, derived from the mainline traffic features, is fed into two parallel modules that share the same architecture but serve different objectives.
@@
Encoding. To enable the model to focus on the spatial dimension, the input embedding $E$ is first transposed along the spatial ($M$) and ramp ($P$) dimensions. The transposed tensor is then fed into the spatial encoder. After encoding and layer normalization, the result is transposed back to the original dimension order to form the spatial representation $H^{(S)}\in\mathbb{R}^{M\times P\times D}$. This process can be expressed as:
$H^{(S)}=\text{LayerNorm}\big(\text{EncoderS}(E^{T})^{T}\big)$ (3)
where $E^{T}\in\mathbb{R}^{M\times P\times D}$denotes the transpose of $E$ along its last two dimensions.
@@
Decoding. The spatial representation $H^{(S)}$ is first passed through a linear mapping layer and then transposed again to match the input requirements of the spatial decoder. The decoder output is normalized and passed through the final output layer, followed by a final transpose to generate the spatially reconstructed sequence $\hat{Y}^{(S)}\in\mathbb{R}^{M\times P\times L}$:
$\hat{Y}^{(S)}=\Big[\text{OutputLayer}\big(\text{LayerNorm}(\text{DecoderS}(\text{Linear}(H^{(S)})^{T}))\big)\Big]^{T}$ (4)
@@
Temporal Reconstruction Module (TAE). This module shares the same architecture as SAE but focuses on capturing long-range temporal dependencies.
Encoding. In the temporal domain, the input embedding $E\in\mathbb{R}^{M\times P\times D}$ is directly fed into the temporal encoder. The output is normalized to obtain the temporal representation $H^{(T)}\in\mathbb{R}^{M\times P\times D}$:
$H^{(T)}=\text{LayerNorm}(\text{EncoderT}(E))$ (5)
Decoding. The temporal representation $H^{(T)}$ is linearly mapped and directly input into the temporal decoder DecoderT without any transposition. The decoder output is normalized and passed through the output layer to obtain the temporally reconstructed sequence $\hat{Y}^{(T)}\in\mathbb{R}^{M\times P\times L}$:
$\hat{Y}^{(T)}=\text{OutputLayer}\big(\text{LayerNorm}(\text{DecoderT}(\text{Linear}(H^{(T)})))\big)$ (6)
The main difference between the two modules lies in the dimension processing strategy. The SAE module reorganizes the spatiotemporal dimensions via a transpose operation, allowing the Transformer to model dependencies across ramps and mainline. In contrast, the T-AE module preserves the original dimension order, focusing on capturing long-range temporal dependencies.
@@
Both modules are optimized by computing reconstruction losses between $\hat{Y}^{(S)}$ and $\hat{Y}^{(T)}$ and the true ramp sequences. SAE captures spatial heterogeneity among ramps, while TAE models long-range temporal patterns. By reconstructing ramp traffic based on mainline flow features, the model implicitly learns the spatiotemporal dependencies between ramps and the mainline. The mainline acts as the traffic backbone, influencing ramp inflows and outflows and reflecting traffic propagation and interaction. Consequently, the reconstructed ramp sequences encode both spatial correlations and temporal dynamics, providing complementary spatiotemporal representations for downstream ramp traffic prediction tasks.

### v2

@@ 2.2 Ramp flow prediction
5 Model Evaluation: Includes Validation Method (VM), Real-Time Capability (RTC), and Scenario Generalization Ability (SGA); RW: Real-World, SB: Simulation/Benchmark; L: Low, M: Medium, H: High.
To alleviate the reliance on training data, transfer learning is introduced into ramp flow estimation. The core idea is to transfer knowledge learned from a data-rich source domain (e.g., mainline) to a data-scarce target domain (e.g., ramps) [zhuang2020comprehensive], providing a feasible solution for ramps without historical data. Although traffic states differ across highway segments, they exhibit inherent correlations, creating a natural scenario for the application of transfer learning [zhuang2020comprehensive]. In ramp flow prediction, transfer learning is mainly applied in two ways: feature-representation transfer, which reduces the discrepancy in marginal distributions between the source and target domains, and parameter transfer, which mitigates differences in conditional distributions. Combining these strategies can significantly improve prediction accuracy and generalization capability in the target domain [zhang2023transfer]. Transfer learning methods encounter new challenges when applied to structurally complex interchange scenarios. First, most prior studies focus on ramps with simple structures. Interchanges, however, have diverse layouts and complex ramp systems [garber1999guidelines]. This makes it difficult to directly apply existing models. As shown in Table 1, they exhibit limited scene generalization ability (SGA) in interchange scenarios. Second, the data environment has evolved. With the deployment of ETC gantries, ramp flow data can now be obtained through vehicle path matching. However, ramp flow prediction faces the ”real-time blind spot” problem in Section 1. Transfer learning methods usually rely only on mainline data and ignore the valuable historical ramp data provided by ETC systems. This limits the ability of models to capture the complex spatiotemporal dependencies between mainline and ramp flows. Therefore, the key challenge is to use limited historical data effectively while handling incomplete mainline inputs. This is essential for developing accurate predictive models for ramps in complex interchanges.
2.3 Pre-training framework for traffic prediction
Inspired by successful practices in the field of natural language processing (NLP), pre-training has become a key paradigm for improving model performance [liu2023pre]. The Transformer architecture, due to its strong ability to handle long sequences, forms the core of such pretrained models [vaswani2017attention]. In this context, researchers extend the idea of pre-training to time series modeling. Through self-supervised learning, models extract effective feature representations from large-scale historical data. This enhances the performance of downstream prediction tasks [shao2022pre, chi2025dynamic, WOS:001347142804014]. Among these methods, MAE [he2022masked] has become a mainstream pre-training strategy. The core task of MAE is to mask part of the observed data and reconstruct it using the remaining visible data. This allows the model to learn deep relationships within the data. Such a property makes MAE naturally suitable for handling missing data, providing a new approach to improving the robustness of traffic prediction models.
Existing pre-training frameworks explore various model structures and masking strategies. Rows 4 to 10 of Table 1 presents a comparison of different pre-training models for traffic flow prediction. TST [zerveas2021transformer] uses a masked Transformer architecture to learn representations of multivariate time series. Other models, including STEP [shao2022pre], DGCN-PTL [chi2025dynamic], and STDMAE [WOS:001347142804014], further integrate the intermediate representations generated by MAE into downstream predictors to improve performance. Notably, STDMAE proposes a spatial-temporal decoupled masking mechanism. It performs masking and reconstruction separately along the temporal and spatial dimensions. This allows the model to better capture spatiotemporal heterogeneity. To model temporal dependencies more precisely, TS-MAE [liu2025ts] constructs a masked autoencoder based on neural ordinary differential equations (Neural ODEs). This enables continuous modeling of evolution patterns. TS-MAE also uses a stepwise masking strategy to enhance learning robustness. In addition, some studies incorporate graph structures into pre-training. PesNet [fang2024prestnet] uses graph convolution operators to recover masked data. JointSTNet [cai2024jointstnet] introduces a spatial graph capsule module and a temporal gating module. It also combines an adaptive masking strategy to improve the extraction and learning of complex spatiotemporal relationships. These models exhibit strong real-time capability and scene generalization ability. Their main difference lies in the duration of the input data.
Although these studies demonstrate the advantages of pre-training in capturing long term and diverse spatiotemporal features, most pre-training tasks focus on self reconstruction using historical data. That is, they reconstruct missing parts of the same source traffic flow. However, it cannot directly solve the problem in interchange scenarios. In these cases, data from one modality (mainline traffic) must be used to predict another modality (ramp traffic). To address this, the research propose a proxy reconstruction pre-training task. The model uses mainline traffic flow data to reconstruct ramp traffic flows. In this framework, the model can learn more intrinsic patterns between mainline and ramp flows. Moreover, masking can be applied to the mainline inputs. This simulates real-world scenarios with partial missing data and enhances the robustness of the prediction model.
In summary, although ramp flow prediction makes great progress, current models still have some limitations. To compare the strengths and weaknesses of different approaches more clearly, this study reviews ramp flow prediction methods together with recent pre-training frameworks. Traditional approaches mainly use traffic flow models and dynamic OD estimation. They can predict ramp flows under certain conditions. However, they require complete and real-time data, and the computation cost is high. Machine learning methods reduce the impact of missing data by using additional static or semi-static features. But they rely heavily on auxiliary information, which limits scalability. Transfer learning methods transfer knowledge from mainline flows to ramps. This improves prediction accuracy, but they remain insufficient in complex interchange scenarios. Pre-training approaches take a different path. They learn representations from historical data through self-supervised learning. These methods show strong robustness and generalization. They also perform well under missing data conditions.
@@ 3 Methodology
This section describes the proposed model in detail. The overall framework of the model is illustrated in Figure 2. The STDAE model consists of two components: a pre-training model and a downstream prediction model. In the pre-training stage, an STDAE structure is designed to more efficiently capture the spatio-temporal dependencies of traffic data. This structure comprises two mutually independent modules: a Temporal AutoEncoder (TAE) and a Spatial AutoEncoder (SAE). Both modules share the same architecture but focus separately on reconstructing features in the temporal and spatial dimensions, respectively, to generate temporal and spatial representations of ramp traffic flow states. In the downstream prediction stage, these representations are processed through a multi-layer perceptron (MLP) module and fused with the hidden layer of the prediction model to obtain the final forecast of ramp traffic flow.
@@ 3.1 Problem Statement
In this study, the short-term prediction of ramp traffic flow at interchanges is defined as a temporal prediction problem. Consider a hub interchange with two mainlines crossing, where the mainlines consist of $N=8$ directions, including four upstream and four downstream directions for the two intersecting lines. There are $M=12$ traffic movements (ramps), with each ramp connecting the upstream and downstream mainline flows from two corresponding directions.
@@
To capture the dependencies among ramps, a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is constructed, where each node corresponds to a ramp and each edge denotes spatial or functional correlations among ramps. The graph has $M$ nodes corresponding to the $M$ ramps, and its adjacency matrix $\mathbf{A}\in\mathbb{R}^{M\times M}$ encodes the pairwise relationships among ramps:
@@
In this study, a fully connected graph by default is assumed, i.e., every ramp node is connected to all other ramp nodes, allowing the model to capture interactions among all ramps. This $M\times M$ fully connected graph structure is critical for accurate ramp traffic flow prediction.
3.2 The pre-training model
During the pre-training stage, a conventional autoencoding scheme that encodes the input into hidden representations and then reconstructs the input itself is not adopted. Instead, the model reconstructs the ramp traffic flow sequences by leveraging the long historical time series of upstream and downstream mainline traffic flows. Spatiotemporal decoupling is the key strategy for modeling the spatiotemporal relationships between mainline and ramp traffic flows. Its core idea is to separate spatial dependencies from temporal dependencies and extract their features independently, thereby capturing complex spatiotemporal dynamics more accurately. In addition, an optional masking module is incorporated to address cases where part of the mainline traffic flow data is missing.
Figure 2: STDAE Pre-training and Prediction Framework
@@ 3.2.1 Input Feature Encoding
The input feature encoding process of the model consists of two core stages: Feature Fusion and Feature Encoding. The goal is to transform raw spatiotemporal data into informative embeddings suitable for STDAE.
The objective of feature fusion is to integrate mainline traffic data to construct a comprehensive feature representation for each ramp. The raw input consists of sequential mainline traffic data, including traffic flow, speed, day of the week, time of day, and other attributes. This data is represented as $\mathbf{V}_{t-({T}_{long}-1):t}\in\mathbb{R}^{{T}_{long}\times N\times F}$, where $T_{\text{long}}$ denotes the length of the input time series. Next, these mainline features are reorganized and aggregated to construct ramp features. For the $m$-th ramp, $m\in\{1,\dots,M\}$, let $U(m)$ and $D(m)$ denote the indices of its upstream and downstream mainline detectors, respectively. At any time step $\tau$, the feature vector of ramp $m$ is formed by concatenation:
$\mathbf{x}_{\tau,m}=\text{Concat}_{n\in U(m)\cup D(m)}\big(\mathbf{v}_{\tau,n}\big),$ (3)
where $\mathbf{v}_{\tau,n}\in\mathbb{R}^{F}$ represents the feature vector of mainline detector $n$ at time $\tau$.
By applying this procedure to all $M$ ramps and $T_{\text{long}}$ time steps, the fused feature tensor $\mathbf{X}\in\mathbb{R}^{T_{\text{long}}\times M\times C}$ is obtained. where $C=(\lvert U(m)\rvert+\lvert D(m)\rvert)\cdot F$ depends on the number of upstream and downstream connections. This step is applied during both pre-training and prediction stages to ensure consistent input representation.
The feature encoding stage transforms the fused tensor $\mathbf{X}$ into the final input embedding $\mathbf{E}$, consisting of masking, patching & embedding, and positional encoding.
To simulate missing mainline data, spatial and temporal masks, denoted as $\mathbf{M}{\mathrm{s}}$ and $\mathbf{M}{\mathrm{t}}$, are defined as binary tensors of the same shape as $\mathbf{X}$, where 1 indicates retained data and 0 indicates masked data.The masking operation is applied as:
$\mathbf{X}^{\prime}=\mathbf{X}\odot\mathbf{M}_{\mathrm{s}}\odot\mathbf{M}_{\mathrm{t}},$ (4)
where $\odot$ denotes the Hadamard product.
For the spatial mask:
[…]
For the temporal mask:
[…]
If the data is complete, no masking is applied, i.e., $\mathbf{X}^{\prime}=\mathbf{X}$.
To reduce computational cost for long sequences $T_{\text{long}}$, the time dimension is divided into $P$ non-overlapping patches, each of length $L=T_{\text{long}}/P$. Let the reshaped input for the convolutional patch embedding layer be $\mathbf{X}_{\text{conv}}\in\mathbb{R}^{M\times C\times T_{\text{long}}}$. The 2D convolution kernel $\mathbf{W}_{\text{conv}}\in\mathbb{R}^{D\times C\times L\times 1}$ with bias $\mathbf{b}_{\text{conv}}\in\mathbb{R}^{D}$ is applied:
$\mathbf{E}_{p}=\text{Conv2D}(\mathbf{X}_{\text{conv}})\in\mathbb{R}^{M\times P\times D}.$ (7)
Before adding positional information, the embedding $\mathbf{E}_{p}$ is normalized along the feature dimension:
$\mathbf{E}_{\mathrm{norm}}=\text{LayerNorm}(\mathbf{E}_{p}),$ (8)
where for any vector $\mathbf{z}\in\mathbb{R}^{D}$:
$\text{LayerNorm}(\mathbf{z})=\frac{\mathbf{z}-\mu}{\sqrt{\sigma^{2}+\epsilon}}\odot\gamma+\beta,$ (9)
with $\mu$ and $\sigma^{2}$ being the mean and variance of $\mathbf{z}$, $\gamma,\beta\in\mathbb{R}^{D}$ learnable parameters, and $\epsilon$ a small constant.
To encode spatial and temporal positions, a 2D sinusoidal positional encoding is defined [wang2021translating]:
$(\mathbf{E}_{\mathrm{pos}})_{m,p,d}=\begin{cases}\sin\left(\dfrac{m}{10000^{2i/D_{\text{half}}}}\right),&d=2i,\;d<D_{\text{half}},\\[8.0pt] \cos\left(\dfrac{m}{10000^{2i/D_{\text{half}}}}\right),&d=2i+1,\;d<D_{\text{half}},\\[8.0pt] \sin\left(\dfrac{p}{10000^{2i/D_{\text{half}}}}\right),&d=D_{\text{half}}+2i,\\[8.0pt] \cos\left(\dfrac{p}{10000^{2i/D_{\text{half}}}}\right),&d=D_{\text{half}}+2i+1,\end{cases}$ (10)
where $D_{\text{half}}=D/2$ and $i$ indexes the feature dimensions.
Finally, the normalized embedding is combined with positional encoding to form the final input:
$\mathbf{E}=\mathbf{E}_{\mathrm{norm}}+\mathbf{E}_{\mathrm{pos}}\in\mathbb{R}^{M\times P\times D}.$ (11)
This tensor $\mathbf{E}$ contains both the semantic information of the original data and the spatiotemporal structure, serving as input to subsequent Transformer-based models.
@@ 3.2.2 STDAE
The embedding $\mathbf{E}$, derived from the mainline traffic features, is fed into two parallel modules that share the same architecture but serve different objectives [WOS:001347142804014].
@@
To enable the model to focus on the spatial dimension, the input embedding $\mathbf{E}$ is first transposed along the spatial ($M$) and ramp ($P$) dimensions. The transposed tensor $\mathbf{E}^{T}\in\mathbb{R}^{M\times P\times D}$ is then fed into the spatial encoder. Within each encoder layer, the spatial self-attention is computed among all ramps at each time step, formulated as:
$\text{Attn}^{(S)}(\mathbf{Q}_{S},\mathbf{K}_{S},\mathbf{V}_{S})=\text{Softmax}\left(\frac{\mathbf{Q}_{S}\mathbf{K}_{S}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}_{S},$ (12)
where $\mathbf{Q}_{S}=\mathbf{E}^{T}\mathbf{W}_{Q}^{(S)}$, $\mathbf{K}_{S}=\mathbf{E}^{T}\mathbf{W}_{K}^{(S)}$, and $\mathbf{V}_{S}=\mathbf{E}^{T}\mathbf{W}_{V}^{(S)}$ denote the query, key, and value matrices, and respectively. This attention mechanism allows each ramp to attend to all other ramps (and the mainline) at the same time step, thereby modeling the spatial correlations of traffic states.
After encoding and layer normalization, the result is transposed back to the original dimension order to form the spatial representation $\mathbf{H}^{(S)}\in\mathbb{R}^{M\times P\times D}$:
$\mathbf{H}^{(S)}=\text{LayerNorm}\big(\text{EncoderS}(\mathbf{E}^{T})^{T}\big).$ (13)
@@
The spatial representation $\mathbf{H}^{(S)}$ is first passed through a linear mapping layer and then transposed again to match the input requirements of the spatial decoder. The decoder reconstructs the spatial sequences through the same type of attention operation:
$\text{Attn}^{(S)}_{\text{dec}}(\mathbf{Q}^{\prime}_{S},\mathbf{K}^{\prime}_{S},\mathbf{V}^{\prime}_{S})=\text{Softmax}\left(\frac{\mathbf{Q}^{\prime}_{S}\mathbf{K}^{\prime\top}_{S}}{\sqrt{d_{k}}}\right)\mathbf{V}^{\prime}_{S},$ (14)
followed by normalization and an output projection to obtain the reconstructed spatial sequence:
$\hat{\mathbf{Y}}^{(S)}=\Big[\text{OutputLayer}\big(\text{LayerNorm}(\text{DecoderS}(\text{Linear}(\mathbf{H}^{(S)})^{T}))\big)\Big]^{T}.$ (15)
@@
Temporal Reconstruction Module (TAE). This module shares the same architecture as SAE but focuses on capturing long-range temporal dependencies rather than spatial ones.
In the temporal domain, the input embedding $\mathbf{E}\in\mathbb{R}^{M\times P\times D}$ is directly fed into the temporal encoder. Here, the temporal self-attention operates along the time dimension, enabling each time step to attend to all other time steps for the same ramp and mainline flow:
$\text{Attn}^{(T)}(\mathbf{Q}_{T},\mathbf{K}_{T},\mathbf{V}_{T})=\text{Softmax}\left(\frac{\mathbf{Q}_{T}\mathbf{K}_{T}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}_{T},$ (16)
where $\mathbf{Q}_{T}=\mathbf{E}\mathbf{W}_{Q}^{(T)}$, $\mathbf{K}_{T}=\mathbf{E}\mathbf{W}_{K}^{(T)}$, and $\mathbf{V}_{T}=\mathbf{E}\mathbf{W}_{V}^{(T)}$. This formulation allows the model to capture long-range temporal dependencies, including recurrent congestion patterns and propagation delays between mainline and ramps. After applying temporal encoding and layer normalization, the temporally encoded representation is obtained as:
$\mathbf{H}^{(T)}=\text{LayerNorm}(\text{EncoderT}(\mathbf{E})).$ (17)
The temporal representation $\mathbf{H}^{(T)}$ is linearly mapped and directly input into the temporal decoder without transposition. The decoder applies the same temporal attention mechanism:
$\text{Attn}^{(T)}_{\text{dec}}(\mathbf{Q}^{\prime}_{T},\mathbf{K}^{\prime}_{T},\mathbf{V}^{\prime}_{T})=\text{Softmax}\left(\frac{\mathbf{Q}^{\prime}_{T}\mathbf{K}^{\prime\top}_{T}}{\sqrt{d_{k}}}\right)\mathbf{V}^{\prime}_{T},$ (18)
followed by layer normalization and the output projection:
$\hat{\mathbf{Y}}^{(T)}=\text{OutputLayer}\big(\text{LayerNorm}(\text{DecoderT}(\text{Linear}(\mathbf{H}^{(T)})))\big).$ (19)
The main difference between SAE and TAE lies in the dimension along which attention is computed. SAE transposes the embedding tensor to allow attention across ramps and mainline flows at a fixed time step, thereby modeling spatial dependencies. In contrast, TAE preserves the original dimension order to perform attention across time steps, thereby modeling temporal dependencies. Formally, SAE focuses on:
[…]
while TAE focuses on:
[…]
@@
Both modules are optimized by computing reconstruction losses between $\hat{\mathbf{Y}}^{(S)}$ and $\hat{\mathbf{Y}}^{(T)}$ and the true ramp sequences. The SAE captures spatial heterogeneity among ramps, while the TAE models long-range temporal patterns. By reconstructing ramp traffic based on mainline flow features, the model implicitly learns the spatiotemporal dependencies between ramps and the mainline. The mainline acts as the traffic backbone, influencing ramp inflows and outflows and reflecting traffic propagation and interaction. Consequently, the reconstructed ramp sequences encode both spatial correlations and temporal dynamics, providing complementary spatiotemporal representations for downstream ramp traffic prediction tasks.